Calculate the exterior angle of a regular polygon with an interior angle of
a) 20°
b) 45°

Found 2 solutions by math_helper, Alan3354:
Answer by math_helper(2461)


a) IF such a polygon existed, the exterior angle would be 180-20 = 160
Note that the number of sides, n, must obey 160 = 180 - (360/n) which has no integer solution for n.
b) Similar

Answer by Alan3354(69443)

Calculate the exterior angle of a regular polygon with an interior angle of
a) 20°
b) 45°
=====================
Not possible.
========================
But,
Calculate the interior angle of a regular polygon with an exterior angle of
a) 20°
Int = 180 - 20 = 160
-------------------
b) 45°
Int = 180 - 45 = 135
----------------------------
Silly problem.